OpenStudy (anonymous):

I'm trying to determine the other angles of a right triangle using a circle. The central angle that shows where the terminal side of the triangle is (3Pi)/2 which I converted to 270 degrees. However, if that is the terminal side of the triangle, that means the only angle I have is 90 degrees. I have no side measurements and no radius for the circle.

OpenStudy (anonymous):

Picture?

OpenStudy (anonymous):

 Homework Directions: For 3-6 sketch the terminal side of the angle. Draw the reference triangle and label the lengths of the sides. Then List all the 6 trig ratios. \[\theta = \frac{ 3\pi }{ 2 }\]